### v2

@@ Abstract
Classical data can be copied and re-used for computation, with adverse consequences economically and in terms of data privacy. Motivated by this, we formulate problems in one-way communication complexity where Alice holds some data and Bob holds $m$ inputs, and he wants to compute $m$ instances of a bipartite relation on Alice’s data and each of his inputs. We call this the asymmetric direct sum question for one-way communication. We give a number of examples where the quantum communication complexity of such problems scales polynomially with $m$, while the classical communication complexity depends at most logarithmically on $m$.
For these examples, data behaves like a consumable resource when the owner stores and transmits it as quantum states. We show an application to a strategic data-selling game, and discuss other potential economic implications.
@@
        Dar Gilboa**footnotemark: *         Siddhartha Jainfootnotemark:         Jarrod McClean
@@ 1 Introduction
As statistical models fitted to large datasets are being usefully applied to problems in various fields of science and engineering [13, 16, 38], the use of proprietary data for training or inference raises concerns of data privacy and adequate compensation for the data owner. The destructive nature of measurement in quantum mechanics has the potential to change this picture. In order to model this scenario we introduce the asymmetric direct sum question in one-way communication complexity. Informally we say a relation $\mathcal{R}$ has an asymmetric direct sum property for communication model $\cal M$ if the communication complexity of computing $\mathcal{R}(x,y_{1})\cdot\mathcal{R}(x,y_{2})\cdot\ldots\cdot\mathcal{R}(x,y_{m})$ is $\Omega(m)$ times the communication complexity of computing $\mathcal{R}(x,y)$. One would expect examples of this in quantum communication when the state Alice sends to Bob undergoes destructive measurement, and may not be copyable. As such, we refer to problems exhibiting this property as consumable data problems [^1] To make our definition robust, we precisely define consumable data probltems to be those where the scaling is polyomial in $m$.. To our knowledge, for communication complexity such a model is studied here for the first time. The work of Hazan & Kushilevitz [26] is superficially similar but the crucial difference is that in our work Alice receives only one input whereas in theirs Alice has $m$ independent instances.
@@
This consumable data property is easily seen to not hold for all relations when $\cal M$ corresponds to deterministic one-way communication. Since Alice’s message depends only on $x$, Bob can copy the message $m$ times and solve each of his instances. On the other hand, we show that for quantum communication there is a scaling of $\widetilde{\Omega}(\sqrt{m})$ for the Hidden Matching problem in Section 5.1. We conjecture this can be improved to $\Omega(m)$ for $m$ not too large. Our results in Sections 5.2 and 5.3 get a scaling of $\widetilde{\Omega}(m)$ for the problems of sampling from a solution of a linear system and estimating the expectation values of two-outcome observables in a multi-party setting, when the parties holding the observables are restricted to using classical communication and performing single-copy measurements of Alice’s messages. For these problems, we also show that classical one-way communication does not exhibit this scaling. Note that the quantum asymmetric direct sum property only holds for relations which have a $\Omega(m)$ separation between quantum and deterministic communication complexity, otherwise the quantum protocol could mimic the deterministic one.
@@
Our results have an interesting economic interpretation, specifically in a setting where a data holder wishes to maximize the payoff of selling data that other parties wish to use for computation, or prevent unauthorized re-use. Models of markets are concerned chiefly with goods that are consumed during the process of economic production, known as rival goods. However, for almost a century it has been recognized that data and information also play a vital role in economic processes [47]. The ability to cheaply replicate data has long been recognized as its chief distinguishing characteristics compared to other economic resources, and this nonrivality has dramatic consequences [8, 45]. It essentially implies that the (albeit idealized) equilibrium known as perfect competition, in which the price of every good on the market is set by its capacity for increasing output, cannot hold once data is involved. In some sense, one cannot “get their money’s worth” when data is traded, unless there is some external enforcement mechanism that sets prices. Such a mechanism may lead to suboptimal resource allocations, and requires trust between the parties involved. Examples of this can be seen in recent proposals for data markets [5, 31]. Nonrivality of data may also disincentivise the creation of novel datasets, which could be of particular concern as the production rate of public high-quality data in certain modalities is far outstripped by the growth rate of training sets for large models [50].
@@
In contrast to the classical picture, the fragile nature of quantum states suggests that classical data encoded in the amplitudes of a quantum state may be destroyed upon use for computation. For this to be the case, one must first show that a problem of interest can be solved with data encoded in this way. In addition, one must argue that the resulting states cannot be replicated in a similar manner to classical data. There is an inherent tension between these two goals, since while no-cloning is trivial for general quantum states [42], this is no longer the case once states are structured. As a simple example, given a computational basis state, it can be measured in the computational basis without disturbing it and subsequently copied, and thus acts analogously to classical data. A less trivial example is that states encoding boolean functions may also be copied in some cases [1]. It is therefore a priori not obvious whether any problems satisfy these competing demands. The nuanced nature of clonability for both the state and the task motivates the formal study of this problem.
@@
The problems we consider are based on ones that exhibit exponential quantum communication advantages. One may wonder whether such an advantage immediately implies the asymmetric direct sum property. In Section 5.3 we show that this is not the case, by considering the problem of observable estimation in a two-party setting [32, 43], and using shadow tomography [2].
@@ 2.1 Destructive measurement as a resource
The idea of using uncloneability of quantum states as a feature has a long history, starting with the seminal work of Weisner [51] that introduced the notion of quantum money. However, the states used in construction of quantum money schemes typically do not encode or transmit useful information and can benefit from the computational power of pseudo-randomness in quantum state [30]. While no-cloning is easy to show for states with little or no structure, this notion becomes more subtle for structured states, and in particular ones that might be useful in performing computation. Aaronson considered the question of uncloneablity of states that encode classical boolean functions, a problem known as quantum software copy-protection [1, 3]. He showed that the presence of structure enables such states to be cloned unless computational assumptions are made, and even then cannot be ruled out for states that encode functions that can be efficiently learned. The setting we consider can be seen as a distributed generalization of this problem. In the simplest case, evaluating the function of interest requires not only a quantum state in the possession of one player (or the equivalent classical description), but also an observable in the possession of another player.
@@ 2.2 Secure Multi-Party Computation
The ability to prevent the re-use of data for computation can in principle be achieved classically using the tools of secure multi-party computation (MPC). The principal objective of line of work is the evaluation of a function $f(x,y)$ where Alice holds $x$ and Bob holds $y$, in a manner that ensures the security of each player’s input and reveals only $f(x,y)$ to both players. There has been extensive work on this problem in various forms since its formulation by Yao [52, 53] (see [21] for a review). Elegant solutions to this problem are known that involve obfuscating (or garbling) the circuit describing $f$ one gate at a time so as to obscure the inputs of each player [23] (which can be achieved using standard cryptographic primitives such as public-key cryptography) or alternatively based on fully homomorphic encryption [9]. Using MPC, the players can run a protocol that enables the evaluation of $f(x,y)$ for a single pair of inputs. However, if Bob wanted to evaluate $f(x,y^{\prime})$ for some $y^{\prime}\neq y$, the validity of any MPC scheme implies that this would generally be impossible without rerunning the protocol. Since this requires the cooperation of both parties, it could allow one party to control the number of times another party can compute functions of their joint inputs.
MPC is incomparable to the consumability of data studied in this work, which relies on the properties of quantum mechanics. MPC has the benefits of being generic and not requiring the constant overheads associated with fault-tolerant quantum computation. However, MPC has a number of drawbacks compared to consumable data, namely (i) it requires multiple rounds of two-way communication [22] whereas our notion of consumable data requires only a single round of one-way communication, (ii) it requires cryptographic assumptions while we give unconditional results and (iii) it requires coordination between parties e.g. in choosing a cryptosystem to use, while our construction requires no such coordination. The best known classical techniques also have overheads associated with them due to the need to encrypt data, which however may not be fundamental.
@@ 3 Preliminaries & Notation
We denote by $D^{\rightarrow}$ deterministic classical one-way communication complexity. $R^{\rightarrow}_{\varepsilon}$ denotes randomized one-way classical communication complexity with error probability at most $\varepsilon$, in which players are allowed to share an unlimited number of public random bits that are independent of their inputs. We similarly define by $Q^{\rightarrow}_{\varepsilon}$ one-way quantum communication complexity with error probability at most $\varepsilon$. In all cases the one-way restriction implies that only Alice is allowed to send messages to Bob (if there are multiple Bobs, they can communicate among themselves and we do not consider this as part of the complexity of the problem). When the error is a nonzero constant (say $1/3$) we omit the subscript. For formal definitions we refer the reader to textbooks by Nisan & Kushilevitz [35], and Lee & Shraibman [36].
@@ 4 Consumable data
A problem $P$ is said to be a consumable data problem if
@@
Definition 4.2 (Nonconsumable data problem).
A problem $P$ is said to be a nonconsumable data problem if
@@
We refer to the quantity appearing in the lower bound in Definition 4.1 as the consumability rate of $P$. In other words, we can say that a problem $P$ is a consumable data problem if its consumability rate is polynomial, and it is nonconsumable otherwise. There is a subtlety in this definition, in the sense that the benefit of consumability arises when Alice chooses to use a particular communication protocol (typically a quantum one over a classical one) but the definition itself does not specify why she would have such a preference. A natural way to introduce a preference is by formulating a strategic game that involves communication. We provide an example in Section 6.2.
Note that any problem involving only deterministic classical communication must be nonconsumable – every Bob can just copy Alice’s message into his own working space. We also show in Section 5.3 that if $P$ corresponds to a decision problem, then even with quantum communication it must be a nonconsumable data problem. This is because the Bobs can apply the Shadow Tomography protocol [2] (unless the Bobs are only allowed classical communication between themselves and limited quantum memory). Nevertheless, consumability can be proved for certain search problems (with many solutions) solved using randomised or quantum communication. There are a few cases where consumability or nonconsumability can be characterized, which we discuss below:
@@
For any relational problem $\mathcal{R}$ and resource $q$, if the protocol is deterministic one way classical communication, $P=(\mathcal{R},D^{\rightarrow},q)$, then $\frac{c(P^{m})}{c(P)}=1$ and the data is nonconsumable.
@@
This can be achieved by learning the distribution of Bob’s output under the randomness of Alice’s message. Using folklore results (see [15]) this can be done using $K\log m$ times the amount of communication, by learning the $K$-outcome distribution up to error $1/m$. ∎
@@
Akin to Lemma 2, we want to give a protocol for for $P^{m}$ using a protocol for $P$ as a subroutine. We do this by relying on the work of Gong & Aaronson [24] who proved that the distribution of $K$-outcome POVMs on $\log N$ qubits can be learned to constant additive error in $\widetilde{O}(K\log^{2}m\log N)$ copies. ∎
All of these lemmas can be generalised to the setting where $\mathcal{P}_{P}$ is a strict subset of one of these sets, which applies to the sampling models.
We note that we restricted our definition of consumable data to a non-interactive setting of one-way communication. We also focus on unconditional results. If we allow interactivity and computational assumptions, then an analogous capabilities can be (conditionally) achieved by classical communication using Secure Multiparty Communication protocols as outlined in Section 2.2. Our results do not contradict such constructions since in a single application of an MPC protocol, Alice’s messages will not allow Bob to solve the problem for any inputs.
@@ 5.1 Hidden Matching
Hidden matching is a famous example of a relation that exhibits an exponential separation between quantum and classical one-way communication complexity [10]. We prove that for the asymmetric direct sum version of the Hidden Matching problem, quantum data behaves as a consumable resource (while classical data does not).
@@
Problem 1 (Hidden Matching [10]).
@@
A tight lower bound shows that classical communication indeed acts like a nonrival good for this problem. While it is known that $R^{\rightarrow}(\mathrm{HM}_{N})=\Omega(\sqrt{N})$ [10], we believe this is the first characterization of the deterministic complexity of the Hidden Matching problem. The results are consistent with Lemma 1.
@@
$R^{\rightarrow}(\mathrm{MHM}_{N,m})=O(\log m)R^{\rightarrow}(\mathrm{HM}_{N})=O(\sqrt{N}\log m)$
@@
There is a quantum algorithm that solves this problem with probability $1$ using $m\log N$ qubits of communication, which is a trivial repetition of the algorithm of [10]:
@@
We also have the following lower bound on the quantum communication required to solve the problem:
@@
Let $\rho_{\mathbf{X}}$ be density matrix corresponding to the message of length $l$ sent by Alice, of dimension $2^{l}$. By Holevo’s theorem, $I(\mathbf{X}:\mathbf{O})\leq l$. We will show that if the Bobs solve $\mathrm{MHM}_{N,m}$ then $I(\mathbf{X}:\mathbf{O})\geq\Omega(\sqrt{m})$. This gives us the required lower bound.
$I(\mathbf{X}:\mathbf{O})=H(\mathbf{O})-H(\mathbf{O}|\mathbf{X})$. Note that $H(\mathbf{O}|\mathbf{X})=0$ since every Bob’s output is deterministic given the input $\mathbf{X}$. Thus, $I(\mathbf{X}:\mathbf{O})=H(\mathbf{O})$. To make this tuple amenable to analysis, we remove dependencies in the output by considering a spanning forest of the graph induced by $V=\cup_{k}\{i_{k},j_{k}\}=\cup_{k}\{i_{k}\}\cup\cup_{k}\{j_{k}\}$. We have that $|V|\geq\sqrt{m}$ since we had a graph with $m$ distinct edges by construction. Therefore, we get a lower bound of $\Omega(\sqrt{m})$ by Lemma 8. ∎
@@
If we have a tree $T$ on $n$ vertices labelled with variables $x_{1}\ldots x_{n}$, then if $x$ is a uniform random string then the set of random variables $P_{T}=\{b_{uv}|(u,v)\in T\}$ where each $b_{uv}=x_{u}\oplus x_{v}$ with probability at least 2/3 has total entropy at least $\Omega(n)$.
@@
We prove this by induction on the height of the tree $T$, say $h$. If $h=0$, then we have only 1 vertex and the set of parities is empty so the entropy is 0. In the inductive step, we assume that for all trees of height $h-1$, the statement is true. Now, consider any tree $T$ on $n$ vertices of height $h$. Let $L$ be the set of leaves of this tree, and set $k=|V(T)\setminus L|$. We know that the subtree of $T$ up to height $h-1$ has total entropy on the set $P_{T,h-1}$ at least $k-1$. For any vertex $v\in L$, let $p(v)$ be its parent. Then since $x_{v}$ is a uniform random bit and $v$ does not appear in any other parities, $H(x_{v}\oplus x_{p(v)}|P_{T,h-1})=1$. Since $b_{uv}=x_{u}\oplus x_{v}$ with probability 2/3, by concavity of entropy we have that $H(b_{uv}|P_{T,h-1})=\Omega(1)$. We now iterate this argument over all leaves, adding the parities at the leaves to the conditioning. We thus prove our claim. ∎
Combining Lemma 6 and Lemma 7 gives for $m\leq N/2$
$\frac{Q^{\rightarrow}(\mathrm{MHM}_{N,m})}{Q^{\rightarrow}(MHM_{N,1})}=\tilde{\Omega}(\sqrt{m}).$ (5.1)
This implies that Multiple Hidden Matchings is a consumable quantum data problem, with consumability rate $\approx\sqrt{m}$. We believe this result can probably be sharpened to $\tilde{\Omega}(m)$ for $m\ll N$. Note that this is not the case classically. The deterministic lower bound in Lemma 4 also illustrates explicitly that the message sent by Alice to solve a single matching, if composed of raw bits of her input, can be immediately re-used to solve the problem for all possible matchings (since $N+1$ bits will contain the end-points of an edge of any possible perfect matching).

### v3

@@ Abstract
Classical data can be copied and re-used for computation, with adverse consequences economically and in terms of data privacy. Motivated by this, we formulate problems in one-way communication complexity where Alice holds some data $x$ and Bob holds $m$ inputs $y_{1},\ldots,y_{m}$. They want to compute $m$ instances of a bipartite relation $R(\mathord{\cdot},\mathord{\cdot})$ on every pair $(x,y_{1}),\ldots,(x,y_{m})$. We call this the asymmetric direct sum question for one-way communication. We give examples where the quantum communication complexity of such problems scales polynomially with $m$, while the classical communication complexity depends at most logarithmically on $m$. Thus, for such problems, data behaves like a consumable resource that is effectively destroyed upon use when the owner stores and transmits it as quantum states, but not when transmitted classically. We show an application to a strategic data-selling game, and discuss other potential economic implications.
@@
        Dar Gilboa**footnotemark: *         Siddhartha Jainfootnotemark:         Jarrod R. McClean
@@ 1 Introduction
As statistical models fitted to large datasets are being usefully applied to problems in various fields of science and engineering [BMR+20, CTJ+21, MBS+23], the use of proprietary data for training or inference raises concerns of data privacy and adequate compensation for the data owner. The destructive nature of measurement in quantum mechanics has the potential to change this picture. In order to model this scenario we introduce the asymmetric direct sum question in one-way communication complexity. Informally we say a relation $\mathcal{R}$ has an asymmetric direct sum property for communication model $\cal M$ if the communication complexity of computing $\mathcal{R}(x,y_{1})\cdot\mathcal{R}(x,y_{2})\cdot\ldots\cdot\mathcal{R}(x,y_{m})$ is $\Omega(m)$ times the communication complexity of computing $\mathcal{R}(x,y)$. One would expect examples of this in quantum communication when the state Alice sends to Bob undergoes destructive measurement, and may not be copyable. As such, we refer to problems exhibiting this property as consumable data problems [^1] To make our definition robust, we precisely define consumable data probltems to be those where the scaling is polyomial in $m$.. To our knowledge, for communication complexity such a model is studied here for the first time. The work of Hazan & Kushilevitz [HK17] is superficially similar but the crucial difference is that in our work Alice receives only one input whereas in theirs Alice has $m$ independent instances.
@@
We provide a number of examples of consumable data problems when using quantum communication. For the asymmetric sum version of the Hidden Matching problem [BYJK08], we show in Section 5.1 that the one-way communication complexity is $\widetilde{\Omega}(\sqrt{m})$. This follows from an information theoretic argument and generalizes to other settings as discussed in Section 7. We conjecture this can be improved to $\Omega(m)$ for $m$ not too large, and we study a modification of the problem where we show the complexity is indeed $\Omega(m)$. Our results in Sections 5.2 and 5.3 prove a scaling of $\widetilde{\Omega}(m)$ for the problems of sampling from a solution of a linear system and estimating the expectation values of two-outcome observables in a multi-party setting, when the parties holding the observables are restricted to using classical communication and performing single-copy measurements of Alice’s messages. For these problems, we also show that classical one-way communication does not exhibit this scaling. More generally, the consumable data property is easily seen to not hold for all relations when $\cal M$ corresponds to deterministic one-way communication. Since Alice’s message depends only on $x$, Bob can copy the message $m$ times and solve each of his instances. Additionally, the quantum asymmetric direct sum property only holds for relations which have a $\Omega(m)$ separation between quantum and deterministic communication complexity, otherwise the quantum protocol could mimic the deterministic one.
@@
Our results have an interesting economic interpretation, specifically in a setting where a data holder wishes to maximize the payoff of selling data that other parties wish to use for computation, or prevent unauthorized re-use. Models of markets are concerned chiefly with goods that are consumed during the process of economic production, known as rival goods. However, for almost a century it has been recognized that data and information also play a vital role in economic processes [Sch42]. The ability to cheaply replicate data has long been recognized as its chief distinguishing characteristics compared to other economic resources, and this nonrivality has dramatic consequences [Arr62, Rom90]. It essentially implies that the (albeit idealized) equilibrium known as perfect competition, in which the price of every good on the market is set by its capacity for increasing output, cannot hold once data is involved. In some sense, one cannot “get their money’s worth” when data is traded, unless there is some external enforcement mechanism that sets prices. Such a mechanism may lead to suboptimal resource allocations, and requires trust between the parties involved. Examples of this can be seen in recent proposals for data markets [ADS19, JT20]. Nonrivality of data may also disincentivise the creation of novel datasets, which could be of particular concern as the production rate of public high-quality data in certain modalities is far outstripped by the growth rate of training sets for large models [VSH+22].
@@
In contrast to the classical picture, the fragile nature of quantum states suggests that classical data encoded in the amplitudes of a quantum state may be destroyed upon use for computation. For this to be the case, one must first show that a problem of interest can be solved with data encoded in this way. In addition, one must argue that the resulting states cannot be replicated in a similar manner to classical data. There is an inherent tension between these two goals, since while no-cloning is trivial for general quantum states [NC10], this is no longer the case once states are structured. As a simple example, given a computational basis state, it can be measured in the computational basis without disturbing it and subsequently copied, and thus acts analogously to classical data. A less trivial example is that states encoding boolean functions may also be copied in some cases [Aar09]. It is therefore a priori not obvious whether any problems satisfy these competing demands. The nuanced nature of clonability for both the state and the task motivates the formal study of this problem.
@@
The problems we consider are based on ones that exhibit an exponential quantum communication advantage. One may wonder whether such an advantage immediately implies the asymmetric direct sum property. In Section 5.3 we show that this is not the case, by considering the problem of observable estimation in a two-party setting [Kre95, Raz99], and using shadow tomography [Aar18].
Comparison to Chatterjee et al. [CKP25]
A recent work [CKP25] also studied a problem similar to the consumability of the Hidden Matching problem. They proved that if $k$ separated Bobs want to compute $k$ instances of Hidden Matching then the min-entropy of their joint distribution must be $\Omega(kn)$ for at least an exponentially small fraction of Alice’s possible inputs. Our lower bound holds for the stronger model where a single Bob holds all $k$ instances and our result holds in the bounded error setting, which they leave as an open problem.
@@ 2.1 Destructive measurement as a resource
The idea of using uncloneability of quantum states as a feature has a long history, starting with the seminal work of Weisner [Wie83] that introduced the notion of quantum money. However, the states used in construction of quantum money schemes typically do not encode or transmit useful information and can benefit from the computational power of pseudo-randomness in quantum state [JLS18]. While no-cloning is easy to show for states with little or no structure, this notion becomes more subtle for structured states, and in particular ones that might be useful in performing computation. Aaronson considered the question of uncloneablity of states that encode classical boolean functions, a problem he called quantum software copy-protection [Aar09, ALL+21]. He showed that the presence of structure enables such states to be cloned unless computational assumptions are made, and even then cannot be ruled out for states that encode functions that can be efficiently learned. The setting we consider can be seen as a distributed generalization of this problem. In the simplest case, evaluating the function of interest requires not only a quantum state in the possession of one player (or the equivalent classical description), but also an observable in the possession of another player.
The line of work which most closely resembles our work conceptually is that of One-Time Programs, introduced by Goldwasser et al. [GKR08]. The quantum analog of these was defined by Broadbent et al. in 2013 and has recently seen a surge of interest [BGS13, BDS23, CLLZ21, GM24, GLR+24]. The main difference in our definition is that it is information theoretic and thus does not rely on any computational assumptions. We note that a natural definition of One-Time Programs is provably impossible in many cases, and recent works weaken the model to make it more generically feasible. We hence make an orthogonal inquiry into what is achievable in a very strong sense, even if only in special cases.
@@ 2.2 Secure Multi-Party Computation
The ability to prevent the re-use of data for computation can in principle be achieved classically using the tools of secure multi-party computation (MPC). The principal objective of line of work is the evaluation of a function $f(x,y)$ where Alice holds $x$ and Bob holds $y$, in a manner that ensures the security of each player’s input and reveals only $f(x,y)$ to both players. There has been extensive work on this problem in various forms since its formulation by Yao [Yao82, Yao86] (see [EKR18] for a review). Elegant solutions to this problem are known that involve obfuscating (or garbling) the circuit describing $f$ one gate at a time so as to obscure the inputs of each player [GMW87] (which can be achieved using standard cryptographic primitives such as public-key cryptography) or alternatively based on fully homomorphic encryption [AJLA+12]. Using MPC, the players can run a protocol that enables the evaluation of $f(x,y)$ for a single pair of inputs. However, if Bob wanted to evaluate $f(x,y^{\prime})$ for some $y^{\prime}\neq y$, the validity of any MPC scheme implies that this would generally be impossible without rerunning the protocol. Since this requires the cooperation of both parties, it could allow one party to control the number of times another party can compute functions of their joint inputs.
MPC is incomparable to the consumability of data studied in this work, which relies on the properties of quantum mechanics. MPC has the benefits of being generic and not requiring the constant overheads associated with fault-tolerant quantum computation. However, MPC has a number of drawbacks compared to consumable data, namely (i) it requires multiple rounds of two-way communication [GIKR01] whereas our notion of consumable data requires only a single round of one-way communication, (ii) it requires cryptographic assumptions while we give unconditional results and (iii) it requires coordination between parties e.g. in choosing a cryptosystem to use, while our construction requires no such coordination. The best known classical techniques also have overheads associated with them due to the need to encrypt data, which however may not be fundamental.
@@ 3 Preliminaries & Notation
We denote by $D^{\rightarrow}$ deterministic classical one-way communication complexity. $R^{\rightarrow}_{\varepsilon}$ denotes randomized one-way classical communication complexity with error probability at most $\varepsilon$, in which players are allowed to share an unlimited number of public random bits that are independent of their inputs. We similarly define by $Q^{\rightarrow}_{\varepsilon}$ one-way quantum communication complexity with error probability at most $\varepsilon$. In all cases the one-way restriction implies that only Alice is allowed to send messages to Bob (if there are multiple Bobs, they can communicate among themselves and we do not consider this as part of the complexity of the problem). When the error is a nonzero constant (say $1/3$) we omit the subscript. For formal definitions we refer the reader to textbooks by Nisan & Kushilevitz [KN96], and Lee & Shraibman [LS09].
@@ 4 Consumable data
A problem $P$ is said to be a consumable data problem if as $m\rightarrow\infty$
@@
Definition 4.2 (Non-consumable data problem).
A problem $P$ is said to be a non-consumable data problem if as $m\rightarrow\infty$
@@
We refer to the quantity appearing in the lower bound in Definition 4.1 as the consumability rate of $P$. In other words, we can say that a problem $P$ is a consumable data problem if its consumability rate is polynomial, and it is non-consumable otherwise. Note that $m$ is bounded by a function of $n$ so we have a slight abuse of notation, but this will not be a concern for the problems we consider. There is a subtlety in this definition, in the sense that the benefit of consumability arises when Alice chooses to use a particular communication protocol (typically a quantum one over a classical one) but the definition itself does not specify why she would have such a preference. A natural way to introduce a preference is by formulating a strategic game that involves communication. We provide an example in Section 6.2.
Note that any problem involving only deterministic classical communication must be non-consumable – every Bob can just copy Alice’s message into his own working space. We also show in Section 5.3 that if $P$ corresponds to a decision problem, then even with quantum communication it must be a non-consumable data problem. This is because the Bobs can apply the Shadow Tomography protocol [Aar18] (unless the Bobs are only allowed classical communication between themselves and limited quantum memory). Nevertheless, consumability can be proved for certain search problems (with many solutions) solved using randomised or quantum communication. There are a few cases where consumability or nonconsumability can be characterized, which we discuss below:
@@
For any relational problem $\mathcal{R}$ and resource $q$, if the protocol is deterministic one way classical communication, $P=(\mathcal{R},D^{\rightarrow},q)$, then $\frac{c(P^{m})}{c(P)}=1$ and the data is non-consumable.
@@
This can be achieved by learning the distribution of Bob’s output under the randomness of Alice’s message. Using folklore results (see [Can20]) this can be done using $K\log m$ times the amount of communication, by learning the $K$-outcome distribution up to error $1/m$. ∎
@@
Akin to Lemma 2, we want to give a protocol for for $P^{m}$ using a protocol for $P$ as a subroutine. We do this by relying on the work of Gong & Aaronson [GA23] who proved that the distribution of $K$-outcome POVMs on $\log N$ qubits can be learned to constant additive error in $\widetilde{O}(K\log^{2}m\log N)$ copies. ∎
All of these lemmas can be generalized to the setting where $\mathcal{P}_{P}$ is a strict subset of one of these sets, which applies to the sampling models. Note that the setting where Alice only sends samples of a quantum state is the setting of shadow tomography [Aar18], so our lower bounds generalize lower bounds for shadow tomography.
We also define a related notion of strongly consumable data. To do this, we define the problem $SP^{m}=(\mathcal{R}^{m},\mathcal{P}_{SP}^{m},q)$ where $\mathcal{P}_{SP}^{m}$ is the set of protocols which solve a $2/3$ fraction of Bob’s $m$ instances successfully with high probability. Note that this makes Bob’s task easier, hence achieving strong consumability is harder from Alice’s perspective.
Definition 4.3 (Strongly consumable data problem).
A problem $P$ is said to be a consumable data problem if
$\frac{c(SP^{m})}{c(P)}=m^{\Omega(1)}$
Note that in the strongly consumable case, all classical protocols are non-consumable.
Lemma 4.
For any relational problem $\mathcal{R}\subseteq\mathcal{X}\times\mathcal{Y}\times\mathcal{O}$, with $|\mathcal{O}|=K$ and resource $q$, if the protocol is randomized one-way classical communication, $P=(\mathcal{R},R^{\rightarrow},q)$, then $\frac{c(SP^{m})}{c(P)}=O(\log m)$ and the data is not strongly consumable.
Despite this classical impossibility, our results in the next section continue to hold for strongly consumable data in the quantum setting.
@@ 5.1 Hidden Matching
Hidden matching is a famous example of a relation that exhibits an exponential separation between quantum and classical one-way communication complexity [BYJK08]. We prove that for the asymmetric direct sum version of the Hidden Matching problem, quantum data behaves as a consumable resource (while classical data does not).
@@
Problem 1 (Hidden Matching [BYJK08]).
@@
A tight lower bound shows that classical communication indeed acts like a nonrival good for this problem. While it is known that $R^{\rightarrow}(\mathrm{HM}_{N})=\Omega(\sqrt{N})$ [BYJK08], we believe this is the first characterization of the deterministic complexity of the Hidden Matching problem. The results are consistent with Lemma 1, and provide an explicit demonstration of the phenomenon where a message large enough to solve a single instance of a relation can be re-used to solve others.
@@
$R^{\rightarrow}(\mathrm{MHM}_{N})=O(\sqrt{N}\log m)$
@@
There is a quantum algorithm that solves this problem with probability $1$ using $m\log N$ qubits of communication, which is a trivial repetition of the algorithm of [BYJK08]:
@@
We also have the following lower bound on the quantum communication required to solve the problem.
@@
Let $\rho_{\mathbf{X}}$ be density matrix corresponding to the message of length $l$ sent by Alice, of dimension $2^{l}$. By Holevo’s theorem, $\operatorname{I}(\mathbf{X}:\mathbf{O})\leq l$. We will show that if the Bobs solve $\mathrm{MHM}_{N,m}$ then $\operatorname{I}(\mathbf{X}:\mathbf{O})\geq\Omega(\sqrt{m})$. This gives us the required lower bound.
$\operatorname{I}(\mathbf{X}:\mathbf{O})=\operatorname{H}(\mathbf{X})-\operatorname{H}(\mathbf{X}|\mathbf{O})=n-\operatorname{H}(\mathbf{X}|\mathbf{O})$. To make the random variable $\mathbf{O}$ amenable to analysis, we remove dependencies in the output by considering a spanning forest of the graph induced by $V=\bigcup_{k}\{i_{k},j_{k}\}=\cup_{k}\{i_{k}\}\bigcup\cup_{k}\{j_{k}\}$. We have that $|V|\geq\Omega(\sqrt{m})$ since we have a graph with $m$ distinct edges by construction. Take a spanning tree (which has size $\Omega(\sqrt{m})$) and call the collection of these random variables $\mathbf{O_{T}}$. We can write $\operatorname{H}(\mathbf{X}|\mathbf{O})=\operatorname{H}(\mathbf{X}|\mathbf{O_{T}}\mathbf{O^{\prime}})\leq\operatorname{H}(\mathbf{X}|\mathbf{O_{T}})$.
Now, note that parities encoded in $\mathbf{O_{T}}$ introduce $k=\Omega(\sqrt{m})$ binary constraints on the random variable $\mathbf{X}$ over the boolean hypercube, each of which reduces the support of the conditional distribution by a factor of $2$. Thus, we have that $\operatorname{H}(\mathbf{X|O_{T}})\leq\log(2^{n-k})=n-k$ which means that $\operatorname{I}(\mathbf{X}:\mathbf{O})\geq k=\Omega(\sqrt{m})$. This gives us the desired lower bound by Holevo’s theorem. ∎
Combining Lemma 7 and Lemma 8 gives for $m\leq N/2$
$\frac{Q^{\rightarrow}(\mathrm{MHM}_{N,m})}{Q^{\rightarrow}(\mathrm{MHM}_{N,1})}=\tilde{\Omega}(\sqrt{m}).$ (5.1)
This implies that Multiple Hidden Matchings is a consumable quantum data problem, with consumability rate $\approx\sqrt{m}$. Note that this is not the case classically. The deterministic lower bound in Lemma 5 also illustrates explicitly that the message sent by Alice to solve a single matching, if composed of raw bits of her input, can be immediately re-used to solve the problem for all possible matchings (since $N+1$ bits will contain the end-points of an edge of any possible perfect matching).
Remark 1.
We note that in the proof of Lemma 8, even if $\Omega(m)$ Bobs outputted an answer, we would have that $|V|\geq\Omega(\sqrt{m})$. Therefore, the lower bound holds in the average case and Multiple Hidden Matchings is a strongly consumable data problem with quantum communication.
Finally, we define a problem with a slightly differrent definition which does not formally fit into the framework of Consumable Data (Definition 4.1). However, it is conceptually similar and we can achieve a tight lower bound for it.
Problem 3 (Hidden Matching with Many Edges ($\mathrm{HMM}_{N,m}$)).
Alice is given a string $x\in\{0,1\}^{n}$. Bob is given a perfect matching $M$ over $[N]$. Their goal is for Bob to ouput $m$ tuples $(i_{k},j_{k},x_{i_{k}}\oplus x_{j_{k}})$ where $(i_{k},j_{k})\in M$ for all $i\in[m]$.
We now show that the communication complexity of this problems scales linearly in $m$.
@@
$Q^{\rightarrow}(\mathrm{HMM}_{N,m})=\Omega(m)$ for $m\leq N$
@@
We define the distribution to be $\mathbf{X}\sim U(\{0,1\}^{n})$ with a single perfect matching $M$. Then define the random variable $\mathbf{O}$ in the same way as in the proof of Lemma 8, but now we are guaranteed that $|V|\geq m$, because we are requiring that Bob outputs distinct edges from a matching. This implies that $I(\mathbf{O}:\mathbf{X})\geq\Omega(m)$, giving us a lower bound of $\Omega(m)$ by Holevo’s theorem. ∎